OpenStudy (anonymous):

solving with elimination method. 4x-6y=3 6x-4y=-3

hero (hero):

4x - 6y = 3 6x - 4y = -3 Okay, so this one isn't exactly reduceable, like the previous one.

hero (hero):

Also the slopes are not the same, so this one definitely has a solution.

hero (hero):

Basically, now it just comes down to which variable you want to eliminate. It doesn't matter which one you choose. As long as you choose one.

OpenStudy (anonymous):

could you explain how we figure what numbers to use to eliminate the variable.

hero (hero):

After you choose a variable, say x perhaps, you'll want to rewrite both equations such that the the coefficients of both for variable x are the same.

OpenStudy (anonymous):

ok

hero (hero):

The co-efficients being the values affront the variable.

hero (hero):

In order to make coefficents 4 and 6 the same in both equations, we have to multiply both equations by an appropriate number.

OpenStudy (anonymous):

so if we chose x we would use -4x and -6x right

hero (hero):

Well not exactly. the coefficients must be the same. for example: 4x + 3y = 8 4x + 2y = 6 Notice that the coefficients of variable x for both equations are the same. This means we would be able to use elimination method to simply subtract the second equation from the first. Doing that would lead us to y = 2 since subtracting both equations would lead to elimination of the x variable.

hero (hero):

Another example is: 7x + 4y = 3 3x - 4y = 7 Notice that this time, we have a situation where the coefficients of y are opposites of each other. So combining both equations would eliminate the y variable leaving us with 10x = 10 which further simplifies to x = 1

hero (hero):

So when we eliminate variables we must manipulate the equations to get these type of forms.

hero (hero):

So anyway, back to this particular problem. Here we have: 4x - 6y = 3 6x - 4y = -3 In this case, we don't have either of the sample cases I've shown above. Instead, we have to manipulate the equations to get what we want. In order to do that we have to do a couple of things: 1st: We have to find the LCM of 6 and 4 2nd: We have to multiply the first equation by the appropriate factor to get the LCM as the coefficient of variable x, then we have to multiply the second equation by the appropriate factor to get the same LCM as the coefficient of x.

hero (hero):

In this case, the LCM of 4 and 6 is 12 since: 4 x 3 = 12 6 x 2 = 12 This means we have to: Multiply the 1st equation by 3 Multiply the 2nd equation by 2

hero (hero):

When we do that, we get: 12x - 18y = 9 12x - 8y = -6

hero (hero):

Now we have what we want, which is a case where the coefficients of x match. In this case (as described above) we want to subtract the second equation from the first to eliminate x

hero (hero):

And upon doing so, we get: -26y = 15 which is the same as 26y = -15

hero (hero):

Unfortunately, we will not get easy integer values for y. It will be a fraction upon solving for y: 26y = -15 y = -15/26

hero (hero):

To find x, we will need to plug the value for y back into the original equation and afterwards, solve for x.

OpenStudy (anonymous):

now to do that do we just redo everything we just did but with y?

hero (hero):

We already found y. We have to find x now and I just explained how to find x.

OpenStudy (anonymous):

ok it makes sense now sorry i meant x :)